# AttentionPredictor: Temporal Pattern Matters for Efficient LLM Inference

### v2

@@ Abstract
With the development of large language models (LLMs), efficient inference through Key-Value (KV) cache compression has attracted considerable attention, especially for long-context generation. To compress the KV cache, recent methods identify critical KV tokens through heuristic ranking with attention scores. However, these methods often struggle to accurately determine critical tokens as they neglect the temporal patterns in attention scores, resulting in a noticeable degradation in LLM performance. To address this challenge, we propose AttentionPredictor, which is the first learning-based critical token identification approach. Specifically, AttentionPredictor learns a lightweight convolution model to capture spatiotemporal patterns and predict the next-token attention score. An appealing feature of AttentionPredictor is that it accurately predicts the attention score while consuming negligible memory. Moreover, we propose a cross-token critical cache prefetching framework that hides the token estimation time overhead to accelerate the decoding stage. By retaining most of the attention information, AttentionPredictor achieves 16$\times$ KV cache compression with comparable LLM performance, significantly outperforming the state-of-the-art. The code is available at https://github.com/MIRALab-USTC/LLM-AttentionPredictor.
@@ 1 Introduction
Large language models (LLM) like OpenAI o1 (OpenAI, 2024) have shown impressive scalability and effectiveness in tackling complex tasks through long chain-of-thought (CoT) reasoning and multi-turn conversations (Minaee et al., 2024; Huang et al., 2024). However, these long-context tasks require LLMs to handle extremely lengthy contexts, presenting computational and memory challenges for LLMs (Zhou et al., 2024). Specifically, the key-value cache (KV cache), which holds the attention keys and values during generation to prevent re-computations, consumes huge GPU memory (Li et al., 2024a). For example, for a model with 7 billion parameters, the parameters consume only 14 GB of memory whereas the KV cache requires around 72 GB with the 128K prompt length (Yang et al., 2024). As the decoding latency and memory footprint scale with the KV cache, it is important to compress the KV cache as the prompt expands.
@@
Figure 1: A comparison of H2O, Quest, and AttentionPredictor for identifying critical tokens in the next step with history attention score. Our learning-based spatiotemporal predictor captures the dynamic attention patterns and accurately predicts next-step attention scores.
@@
Many attention sparsity-based methods are proposed to compress the KV cache in the sequence dimension. Previous works have shown that a small portion of tokens dominate the attention distribution, substantially determining token generation precision (Liu et al., 2023; Zhang et al., 2023; Ge et al., 2024). Therefore, we can dramatically reduce the cache size by computing attention sparsely with only the critical tokens, while maintaining LLM performance. Cache eviction methods (Zhang et al., 2023; Xiao et al., 2024; Li et al., 2024b) use heuristic ranking with attention scores to identify critical keys and values and evict the less relevant ones. However, heuristic scoring methods can only model static patterns and struggle to identify critical tokens accurately, causing these methods to suffer from a degradation in LLM performance. Cache retrieval methods (Tang et al., 2024; Zhang et al., 2024a) identify critical tokens by approximating attention using compressed keys and the current query. Although effective, these methods demand substantial computational resources, and model accuracy tends to degrade as the compression ratio and retrieval granularity increase. In particular, Tang et al. (2024) experiences a sharp 11% accuracy drop when the retrieving page size increases from 16 to 64. Additionally, retrieval-based methods rely on the query token in the current step, limiting their ability to overlap the estimation time overhead with actual LLM inference computation in cache prefetching systems. Thus, it is essential to accurately and rapidly identify critical tokens of the KV cache.
In our paper, we explore the importance of temporal patterns in attention scores for identifying critical KV cache. Previous research has indicated that attention scores exhibit repetitive patterns (Jiang et al., 2024; Ge et al., 2024) and are intrinsic to the LLM. These findings suggest that identifying attention patterns enables the prediction of critical tokens, making cache compression possible while preserving most information as in multi-tier cache management systems (Hashemi et al., 2018; Li et al., 2021). We further observe that attention patterns have temporal characteristics, such as re-access, sequential, and seasonal, illustrated in Figure 2. To capture the attention patterns, we introduce AttentionPredictor, a time-series prediction approach to predict attention scores for critical token identification, as shown in Figure 1. Since temporal attention patterns are dynamic and difficult to capture with existing heuristic methods, our approach is the first learning-based method to address this challenge. We frame the prediction of attention scores as a spatio-temporal forecasting problem, treating token positions as spatial patterns. We then train a convolutional model to predict next-step attention scores based on historical attention time series, enabling us to accurately identify the most critical tokens during LLM decoding. Notably, the lightweight AttentionPredictor allows for accurate predictions with negligible memory consumption. Additionally, we employ distribution error calibration by periodic computing dense attention, and we apply block-wise attention compression to further improve the efficiency of prediction. By accurately identifying critical tokens, AttentionPredictor retains most of the attention information after the KV cache compression.
To further accelerate LLM decoding, we apply AttentionPredictor to our proposed KV cache management framework, a cross-token KV cache prefetching system. Current LLM serving systems offload the KV cache to the CPU to reduce the GPU memory usage of long-context generation, but CPU-GPU transfer latency of KV cache becomes a significant bottleneck. KV cache prefetching presents a solution to hide cache loading time by asynchronously loading the required KV cache in advance. In contrast to the existing cross-layer approach (Lee et al., 2024), our cross-token method can capitalize on longer transfer times, and adapt to more sophisticated and accurate methods for identifying critical tokens. Additionally, the transfer data for each token is better organized, leading to improved I/O utilization.
We evaluated our method on several representative LLMs with varying KV cache budgets. On the widely used LongBench dataset, our approach achieves comparable accuracy with 16× KV cache compression, outperforming the state-of-the-art by 41%. With 32K context, our prefetching framework achieves a 1.4$\times$ speedup of per token decoding latency. In summary, we make the following contribution:
  • Based on our observation of the temporal pattern in the attention score, we propose AttentionPredictor, the first learning-based critical token identification approach.
  • We propose the first cross-token prefetching framework, which effectively mitigates prediction and transfer delays in the LLM decoding stage.
  • Experiments on the long-context datasets demonstrate that our approach achieves comparable accuracy of LLM with 16× KV cache compression.
@@ 2.1 Efficient LLM Inference
Several efforts have optimized pre-trained model inference efficiency from different perspectives. Speculative decoding methods (Cai et al., 2024a; Li et al., 2024c; Sun et al., 2024; Liu et al., 2024a) use smaller draft models to accelerate the auto-regressive decoding process of LLMs. Specifically, these methods employ a draft model to efficiently predict several subsequent tokens as generation results, which are then validated in parallel using the target LLM. In contrast, our approach focuses on predicting the attention scores of the next token, serving as an estimation for KV cache compression. Other methods include model compression (Frantar et al., 2023; Lin et al., 2024) and inference systems (Kwon et al., 2023b; Sheng et al., 2023; Song et al., 2024; Contributors, 2023; Zheng et al., 2024; Ye et al., 2025).
@@ 2.2 KV Cache Compression
Many methods are dedicated to compressing the KV cache while retaining as much information of attention as possible. Many works have found that attention scores are sparse, so sparse computation can be performed on high-score positions by estimating attention scores.
@@
Cache eviction. These methods use heuristic approaches to identify key-value pairs of high importance and evict the less relevant ones. StreamingLLM (Xiao et al., 2024) observes that the earliest tokens have high attention scores during inference, so it only retains the initial and the recent few tokens. H2O (Zhang et al., 2023) accumulates all historical attention scores as an estimation, but suffers from the accumulation error of scores from the first few tokens being too frequent. SnapKV (Li et al., 2024b) accumulates attention scores within a recent window as an estimate and performs one-time filtering during the prefill stage. MInference (Jiang et al., 2024) inductively defines three attention patterns and pre-assigns each head a fixed compression method, determining hyperparameters by attention scores within a recent window and staying static during the decoding stage. All these methods are heuristic-based and statically model attention patterns. They struggle to capture the dynamic temporal patterns within attention scores accurately.
Cache retrieval. These methods aim to retrieve the most critical tokens with approximate attention scores. Quest (Tang et al., 2024) achieves this by using the current query and paged key to approximate the attention score. However, Quest is sensitive to the page size, and the accuracy significantly drops with large page sizes and small budgets. Similarly, PQCache (Zhang et al., 2024a) applies key quantization during the prefilling stage and reuses these quantized keys in decoding to approximate attention. Since these methods require information from the current step, they cannot use asynchronous computation to cover long estimation durations. Other methods involving KV cache quantization (Liu et al., 2024b; Hooper et al., 2024; Zhang et al., 2024b) and KV cache budget allocation (Cai et al., 2024b; Yang et al., 2024; Feng et al., 2024) are orthogonal to our token scoring approach and can be combined with our AttentionPredictor.
@@
KV cache cross-layer prefetching. These methods hide part of the cache transfer time based on offloading the cache to the CPU. However, as the sequence length increases, the transfer time is too long to be hidden. InfiniGen (Lee et al., 2024) combines cache retrieval and prefetching by approximating the attention score for the next layer to load the critical cache. However, the estimation time increases significantly as the sequence grows, and the inference time for a single layer is insufficient to cover this. In contrast, our cross-token prefetching framework can hide longer estimation and transfer time within the per-token inference time.
3 Observation
@@ 3.1 Attention Temporal Patterns
Our research is motivated by the observation that attention exhibits three distinct temporal patterns: re-access, sequential, and seasonal. As shown in Figure 2, the re-access pattern occurs as vertical lines, which shows repeated attention to specific tokens. The sequential pattern is marked by diagonal lines, which show the attention moves sequentially to the next tokens. Meanwhile, the seasonal pattern is characterized by the periodic recurrence of critical tokens, which appears in Figure 2 as alternating regions of high and uniform attention scores. These patterns are consistently observed across different models and datasets.
3.2 Predict Attention by Time Series Methods
Capturing attention patterns enables accurate prediction of subsequent attention, suggesting that attention is inherently predictable. Given that LLMs are autoregressive, the process of token generation can be naturally modeled as a time series. Consequently, we propose to model attention scores as a spatiotemporal sequence, where they form a time series along the inference dimension and a spatial sequence along the token dimension. This frames attention prediction as a time series forecasting task, leveraging temporal prediction techniques to forecast the next-step attention. In contrast, existing approaches rely on static attention modeling, limiting their ability to adapt to dynamic temporal variations.
@@ 3.2 Predict Attention by Time Series Methods
Figure 2: Visualization of three temporal attention patterns. Re-access shows repeated attention to specific tokens. Sequential shows attention progresses toward the next tokens. Seasonal exhibits periodic recurrence as alternating bands of high and uniform attention scores.
3.3 Association Between Attention Patterns and Query
To better understand why attention follows temporal patterns, we investigate the underlying causes. We find that the high continuity between queries plays a central role in shaping attention behavior. In particular, the query sequence exhibits a cosine autocorrelation of 87% at a one-step lag (in App. A), which highlights its strong similarity. High query autocorrelation, as observed in our work, is also reported in Lee et al. (2024). Based on this, we further analyze the correlation between attention in step $i$ and step $i+1$. Focusing on the logits before Softmax, the attention computation at step $i$ is given as $A_{i}=\frac{1}{\sqrt{D}}\mathbf{Q_{i}}\mathbf{K_{i}}^{\top}=\frac{1}{\sqrt{D}}q_{i}k_{1:i}^{\top}$, where $q_{i},k_{i}\in\mathbb{R}^{D}$ represent the query and key vectors of the token $i$, respectively. Assuming the query at step $i+1$ satisfies the relationship $q_{i+1}=q_{i}+\Delta q$ , the attention computation at step $i+1$ can then be expressed as:
$A_{i+1}$ $=\frac{1}{\sqrt{D}}\mathbf{Q_{i+1}}\mathbf{K_{i+1}}^{\top}$ (1)
$=\frac{1}{\sqrt{D}}q_{i+1}k_{1:i+1}^{\top}$
$=\frac{1}{\sqrt{D}}(q_{i}+\Delta q)k_{1:i+1}^{\top}$
$=\frac{1}{\sqrt{D}}(q_{i}k_{1:i+1}^{\top}+\Delta qk_{1:i+1}^{\top})$
Focusing on the values of $A_{i+1}$ in the first $i$ positions,
$A_{i+1}[1:i]$ $=\frac{1}{\sqrt{D}}q_{i}k_{1:i}^{\top}+\frac{1}{\sqrt{D}}\Delta qk_{1:i}^{\top}$ (2)
$=A_{i}+\Delta A$
Therefore the difference between $A_{i}$ and $A_{i+1}$ is dominated by $\Delta q$. Since $q_{i}$ and $q_{i+1}$ have high similarity, $\Delta q$ is relatively small. Consequently, $\Delta A$ is small and $A_{i}\approx A_{i+1}$. Therefore, the critical tokens of adjacent steps are similar, which is suitable for cross-token prefetching.
4 Method
@@
Figure 3: Overview of AttentionPredictor and cross-token prefetching framework. (a) AttentionPredictor formulates the attention history as a spatiotemporal sequence, and predicts the attention at the next step with a pre-trained model. To enhance efficiency, the attention history is updated in a compressed form at each decoding step. (b) The cross-token prefetching framework asynchronously evaluates critical tokens and fetches KV for the next token during the LLM inference, thereby accelerating the decoding stage.
In this section, we introduce AttentionPredictor, the first learning-based method for identifying critical tokens, along with the cross-token prefetch framework for improved cache management. We begin with the problem formulation for attention prediction in Section 4.1, followed by a description of our novel AttentionPredictor in Section 4.2. Finally, Section 4.3 presents a cross-token prefetch framework that efficiently hides both evaluation and cache loading latencies.
4.1 Problem Formulation
In the language model decoding stage, we denote $\mathbf{Q}_{t}\in\mathbb{R}^{1\times d}$, $\mathbf{K}\in\mathbb{R}^{t\times d}$ as the query tensor and key tensor used for generate token $t$, respectively. Specifically, we denote $\mathbf{K}_{i}\in\mathbb{R}^{1\times d}$, where $i\in\{1,2,\dots,t\}$, as the key tensor for token $i$, and $\mathbf{K}=\mathbf{K}_{1:t}$ as the complete key tensor. The attention at step $t$ is calculated as:
$A_{t}=\text{Softmax}\left(\frac{1}{\sqrt{d}}\mathbf{Q}_{t}\mathbf{K}^{\top}\right),A_{t}\in\mathbb{R}^{1\times t}.$ (3)
The sparsity-based KV cache compression seeks to find a subset of keys with budget $B$ that preserves the most important attention values. Specifically, the set of selectable key positions is $\Gamma=\{\{\mathbf{p}\}=\left\{p_{i}\right\}_{i=1}^{B}|p_{i}\in\{1,2,\ldots,t\},p_{i}\neq p_{j},\forall i,j=1,2,\ldots,B\}$. We define the attention recovery rate as:
$R_{rec}=\frac{\sum_{i=0}^{B}{A_{t,p_{i}}}}{||A_{t}||_{1}},$ (4)
which reflects the amount of information preserved after compression. A higher recovery rate $R_{rec}$ indicates less information loss caused by KV cache compression. Therefore, the goal of KV cache compression can be formulated as finding the positions $\mathbf{p}$ that maximize $R_{rec}$, i.e.,
$\underset{\mathbf{p}\in\Gamma}{\max}\,R_{rec}.$ (5)
To determine the positions $\mathbf{p}$, existing methods typically employ heuristic approaches to score the attention at step $t$, represented as $S_{t}\in\mathbb{R}^{1\times t}$, and then select the top $B$ positions. For example, the well-known method H2O (Zhang et al., 2023) accumulates historical attention scores, where $S_{t}=\sum_{n=1}^{t-1}{A_{n}}$. In this paper, we predict the attention of step $t$ as $\hat{A_{t}}$ and use it as $S_{t}$.
After identifying the critical token positions $\mathbf{p}$, the attention is computed sparsely $A^{\text{sparse}}=\text{Softmax}\left(\frac{1}{\sqrt{d}}\mathbf{Q}{\mathbf{K}^{\text{sparse}}}^{\top}\right)$, with selected keys $\mathbf{K}^{\text{sparse}}=\text{concate}\{\mathbf{K}_{p_{i}}\}$.

### v3

# AttentionPredictor: Temporal Patterns Matter for KV Cache Compression

@@ Abstract
[…]
@@ 1 Introduction
Large language models (LLM) like OpenAI o1 [1] have shown impressive scalability and effectiveness in tackling complex tasks through long chain-of-thought (CoT) reasoning and multi-turn conversations [2, 3, 4, 5]. However, these long-context tasks require LLMs to handle extremely lengthy contexts, presenting computational and memory challenges for LLMs [6]. Specifically, the key-value cache (KV cache), which holds the attention keys and values during generation to prevent re-computations, consumes huge GPU memory [7]. For example, for a model with 7 billion parameters, the parameters consume only 14 GB of memory whereas the KV cache requires around 72 GB with the 128K prompt length [8]. As the decoding latency and memory footprint scale with the KV cache, it is important to compress the KV cache as the prompt expands.
@@
Figure 1: A comparison of H2O, Quest, SeerAttention, and AttentionPredictor for identifying critical tokens in the next step with history attention score. Our learning-based spatiotemporal predictor captures the dynamic attention patterns and accurately predicts next-step attention scores.
@@
Many attention sparsity-based methods have been proposed to compress the KV cache in the sequence dimension. Previous works have shown that a small portion of tokens dominate the attention distribution, substantially determining token generation precision [9, 10, 11]. Therefore, we can dramatically reduce the cache size by computing attention sparsely with only the critical tokens, while maintaining LLM performance. Cache eviction methods [11, 12, 13] use heuristic ranking with attention scores to identify critical keys and evict the less relevant ones. However, heuristic scoring methods can only model fixed patterns and struggle to identify critical tokens accurately, causing these methods to suffer from LLM performance degradation. Recently, SeerAttention [14] and Attention-Gate [15] use learnable modules to model dynamic patterns and retrieve critical tokens. However, they only represent keys or hidden states, rather than directly modeling the attention score distribution. As a result, they often exhibit limited accuracy after compressing KV cache (Table 8 for details). This motivates the need for a more effective solution that directly targets the prediction of attention scores for enhanced KV cache compression.
In our paper, we thus explore the importance of temporal patterns in attention scores for identifying critical KV cache. Previous research indicates that attention scores exhibit repetitive patterns intrinsic to LLMs [10, 16, 17, 18]. We further observe that attention patterns have temporal characteristics, such as re-access, sequential, and seasonal, as illustrated in Figure 2. Our theoretical analysis in Section 3.3 shows these patterns originate from intrinsic model properties such as query similarity and position encoding. These patterns are both stable and predictable across time, forming a two-dimensional (2D) temporal map with local continuity and translation-invariant characteristics.
KV cache compression with critical token identification can be formulated to attention score prediction which is a 2D time series prediction problem. Inspired by the theoretical understanding of attention mechanisms, we introduce AttentionPredictor, a time-series prediction approach to predict attention scores for critical token identification, as shown in Figure 1. Since temporal attention patterns are dynamic and difficult to capture with existing heuristic methods, we learn lightweight but accurate prediction models to address this challenge. Specifically, we design a convolutional module to capture 2D patterns in the input attention scores, allowing a single predictor to be shared across all transformer layers. These choices shrink our unified prediction model to roughly one millionth the size of the LLM (LLaMA-3.1-8B), resulting in negligible memory consumption. In contrast, SeerAttention’s per‑layer prediction model amounts to 101MB in total, whereas ours is only 21KB—just 0.02% of its size. Additionally, we employ distribution error calibration by periodically computing dense attention, and apply block-wise attention compression to further improve the efficiency of prediction. By accurately identifying critical tokens, AttentionPredictor retains most of the attention information after KV cache compression. To further accelerate LLM decoding, we apply AttentionPredictor to our proposed KV cache management framework, a cross-token KV cache prefetching system. In contrast to the existing cross-layer approach [19], our cross-token method can capitalize on longer transfer times, and adapt to more sophisticated and accurate methods for identifying critical tokens.
[…]
@@ 2.1 Efficient LLM Inference
Several efforts have optimized pre-trained model inference efficiency from different perspectives. Speculative decoding methods [20, 21, 22, 23, 24] use smaller draft models to predict multiple future tokens for parallel validation by a target LLM. In contrast, our approach instead predicts the next token’s attention scores to estimate KV cache compression. Inference systems [25, 26, 27, 28, 29] accelerate the prefilling stage with techniques like prefix cache. AttentionPredictor is compatible with them, for it is applied to the decoding stage. At the same time, our block-based sparsification can also be applied to paged attention [25] and other computational accelerations.
@@ 2.2 KV Cache Compression
Many methods compress the KV cache by leveraging the finding that attention scores are sparse, which allows for estimated sparse computation on high-score positions.
@@
Cache eviction. These methods use heuristic approaches to identify key-value pairs of high importance and evict the less relevant ones. StreamingLLM [12] observes that the earliest tokens have high attention scores during inference, so it only retains the initial and the recent few tokens. H2O [11] accumulates all historical attention scores as an estimation, but suffers from the accumulation error of scores from the first few tokens being too frequent. SnapKV [13] accumulates attention scores within a recent window as an estimate and performs one-time filtering during the prefill stage. MInference [16] and FlexPrefill [30] inductively define several attention patterns and determine the compression strategy for each head to accelerate the prefilling stage. All these methods are heuristic-based and statistically model attention patterns. They struggle to capture the dynamic temporal patterns within attention scores accurately.
Cache retrieval. These methods, such as Quest [31], InfLLM [32], and PQCache [33], retrieve critical tokens by estimating attention weights over compressed key representations. However, such compression can degrade fidelity and introduce retrieval errors. In particular, Quest is sensitive to the page size, and the accuracy significantly drops with large page sizes and small budgets, as shown in Figure 8. Instead, our method employs a lightweight temporal predictor to estimate attention weights directly, yielding higher retrieval accuracy. Other methods involving KV cache quantization [18, 34, 35, 36] and KV cache budget allocation [8, 37, 38, 39] are orthogonal to our token scoring approach and can be combined with our AttentionPredictor.
Training-based Approaches. Recently, a number of training-based cache compression approaches have been proposed. MoBA [40] extends the cache retrieval method by integrating sparse attention during training. Similarly, SeerAttention [14] and NSA [41] both train linear models to encode and retrieve key blocks more accurately. However, these approaches typically require training a separate model for each layer, which limits their scalability and generalization. In contrast, our framework employs a single plug-and-play module that can be applied uniformly to all layers and heads within the same LLM. Our model is only 0.02% the size of SeerAttention. Moreover, whereas NSA relies on the fine-tuning of LLMs to achieve optimal results, our method achieves comparable improvements without any additional fine-tuning, making it both more efficient and broadly applicable.
@@
KV cache cross-layer prefetching. These methods hide part of the cache transfer time based on offloading the cache to the CPU. However, as the sequence length increases, the transfer time is too long to be hidden. InfiniGen [19] combines cache retrieval and prefetching by approximating the attention score for the next layer to load the critical cache. However, the estimation time increases significantly as the sequence grows, and the inference time for a single layer is insufficient to cover this. In contrast, our cross-token prefetching framework can hide longer estimation and transfer time within the per-token inference time.
3 Method
In this section, we introduce AttentionPredictor, the first learning-based method to directly predict attention patterns for KV cache compression and attention sparsity., along with the cross-token prefetch framework for improved cache management. We begin with the problem formulation for attention prediction in Section 3.2, followed by the observation and theoretical analysis of temporal attention patterns in Section 3.3. Then we introduce our novel AttentionPredictor in Section 3.4. Finally, Section 3.5 presents a cross-token prefetch framework that efficiently hides both evaluation and cache loading latencies.
3.1 Preliminary
In the language model decoding stage, we denote $\mathbf{Q}_{t}\in\mathbb{R}^{1\times d}$, $\mathbf{K}\in\mathbb{R}^{t\times d}$ as the query tensor and key tensor used for generate token $t$, respectively. Specifically, we denote $\mathbf{K}_{i}\in\mathbb{R}^{1\times d}$, where $i\in\{1,2,\dots,t\}$, as the key tensor for token $i$, and $\mathbf{K}=\mathbf{K}_{1:t}$ as the complete key tensor. The attention score at step $t$ is calculated as: $A_{t}=\text{Softmax}\left(\frac{1}{\sqrt{d}}\mathbf{Q}_{t}\mathbf{K}^{\top}\right)=[a_{t,1},a_{t,2},\dots,a_{t,t}]\in\mathbb{R}^{1\times t}.$
The sparsity-based KV cache compression seeks to find a subset of keys with budget $B$ that preserves the most important attention values. Specifically, the subset of key indices is $S=\{s_{1},s_{2},...,s_{B}\}$, where each $s_{j}$ is an index within the total $t$ tokens. We define the attention recovery rate as:
[…]
which reflects the amount of information preserved after compression. A higher recovery rate $R_{rec}$ indicates less information loss caused by KV cache compression. Therefore, the goal of KV cache compression can be formulated as finding the index set $S$ that maximizes $R_{rec}$.
3.2 Problem Formulation
Accurate next-step attention score modeling is critical to maintain high attention recovery rate with limited KV tokens and improve final LLM generation performance with KV cache compression. Therefore, KV Cache compression with token selection can be formulated as attention score prediction, which is a two-dimensional (2D) time series prediction problem. Specifically, we predict the attention in step $t+1$ as $\hat{A}_{t+1}$ and select the critical token positions with $S=TopB(\hat{A}_{t+1})$. Since stacking attention vectors over time yields a 2D spatiotemporal attention map $\mathcal{A}_{t}=\langle A_{i}\rangle_{i=1}^{t}$, where the temporal axis corresponds to decoding steps and the spatial axis corresponds to token positions in context as Figure 1. We can thus frame attention score prediction as a time series prediction problem. The input to our predictor is a 2D spatiotemporal sequence $\mathcal{A}_{H}=\langle A^{comp}_{i}\rangle_{i=t-H+1}^{t}$, representing the $H$ most recent block-wise compressed attention scores, where $A_{i}^{comp}$ is generated via max-pooling on the original attention score $A_{i}$ (detailed in Sec. 3.4). We then train a lightweight model to predict the attention scores for step $t+1$ as $\hat{A}_{t+1}=\mathcal{F}(\mathcal{A}_{H})$, where $\mathcal{F}(\cdot)$ denotes the model function.
@@ 3.3 Attention Temporal Patterns
Figure 2: Visualization of three predictable temporal attention patterns. Re-access shows repeated attention to specific tokens. Sequential shows attention progresses toward the next tokens. Seasonal exhibits periodic recurrence as alternating bands of high and uniform attention scores.
Observation. We observe that attention exhibits consistent and structured temporal behaviors, which we term attention patterns. At a local level, these patterns are remarkably stable. We have identified three fundamental patterns that reflect strong invariances under spatiotemporal shifts. However, these patterns reveal a complex dynamic globally, exhibiting long-term evolution. To fully understand this dynamic, we will first provide an analysis of the stable patterns.
Three Attention Patterns. Our work is motivated by the observation that attention exhibits three distinct temporal patterns—re-access, sequential, and seasonal, as shown in Figure 2. These patterns collectively suggest that attention evolves with both temporal continuity and structural regularity across tokens. We unify the three local spatiotemporal attention patterns to the following approximate invariance under joint temporal and spatial shifts:
$a_{t,i}\approx a_{t+\delta_{t},i+\delta_{i}},$ (2)
where and $\delta_{t}$ and $\delta_{i}$ represent temporal and spatial offsets, respectively. (1) The re-access pattern corresponds to $(\delta_{t}=1,\delta_{i}=0)$, where the model repeatedly attends to the same tokens. (2) The sequential pattern corresponds to $(\delta_{t}=1,\delta_{i}=1)$, reflecting progressive token-wise attention. (3) The seasonal pattern corresponds to $(\delta_{t}>1,\delta_{i}=0)$, indicating periodic recurrence of attention to fixed positions. While our re-access and sequential patterns resemble shapes defined in MInference [16], our time-series analysis also reveals a novel seasonal pattern, and our dynamic predictor improves prediction accuracy by 5% over fixed-template approaches (see Appendix C.10).
To better understand why attention follows spatiotemporal patterns, we investigate the underlying theoretical mechanisms. We find out that the high continuity between queries and position embedding plays a central role in shaping attention behavior.
High Query Self-similarity. We observe that the query embedding sequence $\{q_{i}\}_{i=1}^{t}$ exhibits strong temporal self-similarity, which is consistent with prior findings in [19]. Specifically, the one-step cosine autocorrelation is 0.87 as detailed in Appendix A.1, indicating a high degree of alignment between queries. As only the most recent KV cache is updated, high query self-similarity indicates high stability of the attention score over consecutive time steps, which is highly predictable.
For theoretical clarity, we define the raw attention scores as $\widetilde{A}_{t}=q_{t}k_{1:t}^{\top}$ omitting both the $1/\sqrt{d}$ factor and the softmax. This allows us to focus on the underlying structure of attention logits, as softmax is a monotonic function and does not alter the relative ranking of attention weights. For step $t+1$, supposing the query vector evolves as $q_{t+1}=q_{t}+\Delta q$ , the attention at step $t+1$ is:
$\widetilde{A}_{t+1}$ $=q_{t+1}k_{1:t+1}^{\top}=(q_{t}+\Delta q)[k_{1:t}\,|\,k_{t+1}]^{\top}=q_{t}k_{1:t}^{\top}+\Delta qk_{1:t}^{\top}+(q_{t}+\Delta q)k_{t+1}^{\top}$ (3)
Focusing on the first $t$ entries of $\widetilde{A}_{t+1}$, we obtain:
$\widetilde{A}_{t+1}[1:t]$ $=q_{t}k_{1:t}^{\top}+\Delta qk_{1:t}^{\top}=\widetilde{A}_{t}+\Delta\widetilde{A}$, (4)
where $\Delta\widetilde{A}=\Delta qk_{1:t}^{\top}$. Noting that $\|\Delta q\|$ is small due to the high autocorrelation, and the key matrix $k_{1:t}$ remains relatively stable, the term $\Delta\widetilde{A}$ is of small magnitude in norm:
$\|\Delta\widetilde{A}\|\leq\|\Delta q\|\cdot\|\mathbf{K}_{i}\|$
Therefore, $\widetilde{A}_{t}\approx\widetilde{A}_{t+1}$ holds pointwise to a good approximation when the query exhibits high self-similarity. This implies that the relative ranking and distribution of attention weights over past tokens change slowly across decoding steps. In other words, attention exhibits temporal smoothness and inter-step consistency, which form the basis of the observed re-access and sequential patterns.
Position Embedding. The attention patterns are closely related to position embedding, particularly Rotary Positional Embedding (RoPE), which applies position-dependent rotations to queries and keys. In RoPE, the query and key vectors are partitioned into $d/2$ groups along dimensions and apply rotational matrices with different frequencies $\theta$ to each group. The attention weights are calculated with RoPE encoding as [17, 42]:
$\text{RoPE}(q_{i})^{\top}\text{RoPE}(k_{j})$ $=\sum_{m=1}^{d/2}\langle q_{i}^{(m)},R((j-i)\theta_{m})\,k_{j}^{(m)}\rangle$ (5)
$=\sum_{m=1}^{d/2}\|q_{i}^{(m)}\|\,\|k_{j}^{(m)}\|\cos\left(\phi^{(m)}+(j-i)\theta_{m}\right),$
where $q_{i}^{(m)},k_{j}^{(m)}\in\mathbb{R}^{2}$ denote the components of $q_{i}$ and $k_{j}$ in the $m$-th group, $\theta_{m}$ is the frequency parameter, and $\phi^{(m)}$ is the angle between $q_{i}^{(m)}$ and $k_{j}^{(m)}$. For the query and key are stable across steps, the formulation reveals that the inner product depends primarily on the relative position jij-i. When $j-i=\Delta$ is fixed, the cosine terms remain stable across decoding steps, producing consistent attention along diagonals, showing as the sequential pattern. Furthermore, the periodicity of the cosine function explains the periodic slashes of sequential patterns observed in Figure 2.
Dynamic Patterns. These foundational patterns are not static and exhibit dynamic evolution even within a single generation sequence. As the decoding process progresses, patterns can shift, appear, or fade as illustrated in Figure 2. This dynamism is driven by the decay of query-key similarity from inherent query shifting, and is especially evident in long-output tasks like reasoning.
Predict Attention with Time Series Methods. Since the query similarity and position embedding are inherent features of LLMs and independent of input, the resulting attention pattern is also inherent to LLMs. This allows a single shared predictor that can generalize across heads and datasets. For the pattern also changes as the sequence grows, this dynamic requires the predictor not only to recover the stable structures but also to adapt to shifts. Building on theoretical insights into attention, we propose that the pattern can be treated as a two dimensional spatiotemporal process suited to time series methods. In the next section, we introduce an architecture that uses this view to capture and forecast the evolution of attention over time.
@@
Figure 3: Overview of AttentionPredictor and cross-token prefetching framework. (a) AttentionPredictor formulates the history attention scores as a spatiotemporal sequence, and predicts the attention at the next step with a pre-trained model. To enhance efficiency, the attention history is updated in a compressed form at each decoding step. (b) The cross-token prefetching framework asynchronously evaluates critical tokens and fetches KV for the next token during the LLM inference, thereby accelerating the decoding stage.